# Towards Real-time Adaptive Anisotropic Image-to-mesh Conversion for Vascular Flow Simulations

### v1

@@ 5.2 Quantitative and Qualitative Analysis
Although the methods utilized in the proposed approach have been shown to satisfy the I2M requirements using the small test cases in our evaluation, additional experimentation (and fine-tuning of each method’s parameters) is required to process larger test cases (such as the complete vascular structure in LABEL:fig:artery2_original_image_1). The mesh deformation scheme employed by our I2M approach produces meshes with a smoothness of $C^{0}$. This smoothness is sufficient for aneurysm geometries such as those in our evaluation [27]; however, future work is required to improve the smoothness of geometries reconstructed from micro-CT images (i.e., some stent geometries). Our approach must be updated to provide a smoothness of $C^{1}$. Given CDT3D’s ability to process CAD models (such as in fig. 2), this approach remains to be tested within a CFD vascular flow simulation that includes CAD models of stents (testing potential avenues of treatment for other aneurysm cases).
@@
Eventually, the tools utilized in the pipeline of the proposed approach should be integrated into a single software. It should be noted that conversions are required between the input and output files of the methods utilized. For example, several of the methods use the VTK library [6]. Because they were each developed at different time periods, one method produces output utilizing a newer version of the library while another (which utilizes an older version) needs to read this as input. Although the conversion times for the small test cases in our evaluation are negligible, this discrepancy between the methods will likely negatively impact performance when the pipeline is tested on large cases. Integrating these methods into a single software will not only eliminate this problem, but will remove the complexity of compiling all codes separately and running them one after the other in the pipeline. Rather than separating image-to-mesh surface adaptation from anisotropic volume generation (and freezing the surface during this step), it would be more effective to perform both simultaneously when converting the segmented image (similar to the isotropic approach in [31]). If integrated with a boundary layer generation method, this can help prevent the inhibition of mesh quality improvement (such as that shown for the second aneurysm case in LABEL:fig:arteries_CDT3D_MR_EL). Finally, the sequential methods must be parallelized (or replaced with other parallel methods) in order to maximize potential performance, especially for large cases.
7 Conclusion.
Presented is a path towards a fast and robust adaptive anisotropic mesh generation method that is designed to help streamline the discretization of complex vascular geometries and CAD-based stent models within the CFD modeling process. The proposed approach combines multiple software tools into a single pipeline to provide the following: (1) image-to-mesh conversion which satisfies quality, fidelity, and smoothness requirements, (2) the generation of a boundary layer grid over the high fidelity surface, (3) a parallel adaptive anisotropic meshing procedure which satisfies the real-time requirement, and (4) robustness, which is satisfied by the pipeline’s ability to process segmented images and CAD models. This pipeline is tested on a segmented image of a carotid cavernous aneurysm. Additionally, the parallel meshing method is also tested using a surface mesh of a middle cerebral artery bifurcation aneurysm. Both aneurysm cases were obtained from the Cerebrovascular Center for Research at Stony Brook University under Institutional Review Board approval. The pipeline is shown to satisfy all the aforementioned requirements. More specifically, the parallel mesh generation method successfully produces high quality isotropic and anisotropic meshes for the aneurysm cases in real-time. Having shown the feasibility of this approach with regards to these requirements, the next steps are to fully parallelize the remaining components of the pipeline to maximize potential performance and to test its integration within a CFD vascular flow simulation. Just as the parallel anisotropic adaptation procedure was tested within aerospace CFD simulations using CAD models (such as that shown in fig. 2 [64]), the method is expected to provide accurate results for CFD vascular flow simulations in real-time when executed on multicore cc-NUMA architectures.
Acknowledgements
@@
This research was sponsored by the Richard T. Cheng Endowment at Old Dominion University, SURA grant No. C2024-FEMT-011-01, and the National Institute of General Medical Sciences of the National Institutes of Health under Award Number 1T32GM140911-03. This study was also partially supported by a pilot award from the Long Island Network for Clinical and Translational Science. The content is solely the authors’ responsibility and does not necessarily represent the official views of the National Institutes of Health. Additionally, this work was supported by the research computing clusters at Old Dominion University. The authors would like to thank Dr. Christos Tsolakis for helping to assemble the tools utilized in the adaptive pipeline of our approach. We’d also like to thank Dr. David Marcum for sharing the AFLR software utilized in this study, and Evangelia Frastali for helping to determine the best parameters for each method tested in our evaluation.

### v3

@@ 7 Future Work
The mesh deformation scheme employed by CBC3D in our I2M pipeline produces meshes with a smoothness of $C^{0}$. This smoothness is sufficient for aneurysm geometries such as those in our evaluation [EwCCBC3D]; however, future work is required to improve the smoothness of geometries reconstructed from micro-CT images (i.e., some stent geometries). Our approach must be updated to provide a smoothness of $C^{1}$. Given CDT3D’s ability to process CAD models, this approach remains to be tested within a medical numerical simulation that includes CAD models of stents (testing potential avenues of treatment for other aneurysm cases).
@@
Eventually, the tools utilized in the pipeline of the proposed approach should be integrated into a single software. It should be noted that conversions are required between the input and output files of the methods utilized. For example, several of the methods use the VTK library [VTKWebsite]. Because they were each developed at different time periods, one method produces output utilizing a newer version of the library while another (which utilizes an older version) needs to read this as input. Although the conversion times for the small test cases in our evaluation are negligible, this discrepancy between the methods will likely negatively impact performance when the pipeline is tested on large cases. Integrating these methods into a single software will not only eliminate this problem, but will remove the complexity of compiling all codes separately and running them one after the other in the pipeline. Rather than separating image-to-mesh surface adaptation from anisotropic volume generation (and freezing the surface during this step), it would be more effective to perform both simultaneously when converting the segmented image (similar to PODM’s approach [foteinos2014high]). If integrated with a boundary layer generation method, this can help prevent the inhibition of mesh quality improvement (such as that shown for the second aneurysm case in Figure 16(a)). Finally, the sequential methods must be parallelized (or replaced with other parallel methods) in order to maximize potential performance, especially for large cases. As shown in [GarnerThesis], parallelizing AFLR is a non-trivial task (and similar challenges would be expected if one wished to generate boundary layer grids in parallel using the method). Parallelizing the sequential components of the pipeline is outside the scope of this paper however and should be addressed in future work.
Most importantly, the next step is to integrate both proposed methods into a numerical simulation. Such a study will focus on comparing the accuracy and error of critical simulation metrics (such as Wall Shear Stress and velocity fields) to determine if the PODM-generated isotropic meshes are sufficient for clinical-grade predictions, or if the CDT3D-generated anisotropic meshes are necessary to accurately predict an aneurysm’s potential rupture. We will also measure the performance of each method within the context of the simulation’s end-to-end execution time to validate their near real-time performance capabilities. This future work is essential to confirm the end-user productivity of these approaches.
8 Conclusion
Presented are two techniques that are designed to help streamline the discretization of complex vascular geometries within the numerical modeling process. The first approach combines multiple software tools into a single pipeline to generate an adaptive anisotropic mesh from a segmented image of a carotid cavernous aneurysm, providing near real-time performance, good quality, fidelity, smoothness and robustness. The adaptive anisotropic mesh generation method is also tested using a surface mesh of a middle cerebral artery bifurcation aneurysm. The anisotropic method leverages a hierarchical load balancing model while also using an optimized local reconnection algorithm that is three times faster than its previous implementation from past studies [EwCCDT3D]. The second presented technique focuses on generating adaptive isotropic meshes while utilizing a user-defined sizing function, also providing real-time performance, good quality, and fidelity. When utilizing 96 CPU cores on a single, multicore node on Purdue University’s Anvil supercomputer [Anvil2022], the adaptive isotropic method (PODM) generates about 50 million elements in less than a minute while the adaptive anisotropic method (CDT3D) generates approximately the same amount in about 5 minutes. Such performance encourages an investment into a multi-core, shared memory machine that can potentially be utilized within a clinical setting (e.g., a single AMD processor with up to 96 CPU cores can cost about \$10K). Alternatively, end users may benefit from leveraging the concurrency offered by a virtual machine (such as Google’s Compute Engine [GoogleComputeEngine]) when integrating the presented methods into their numerical simulation workflows.
Both techniques (converting medical images to either isotropic or anisotropic meshes) are presented as feasible options for end users, depending on the simulation and accuracy (i.e., tradeoff between speed and the mesh type) needed. For example, isotropic meshes of cerebral vasculature have been utilized within hemodynamic simulations to accurately predict flow conditions of stents post-deployment [LargeMeshesStents2019]. This study notes the expensive computational cost of generating meshes with up to 200 million elements, which could potentially be generated quickly using a method like PODM. On the other hand, the study in [AnisotropicCerebralAneurysms2013] shows that anisotropic meshes may offer more accurate results when including directional information (important for metrics like blood flow velocity). Just as the parallel anisotropic adaptation procedure (CDT3D) was tested within aerospace CFD simulations using CAD models [EwCCDT3D], the method is expected to provide accurate results for medical numerical simulations involving CAD-based stent models in near real-time when executed on large multicore cc-NUMA (shared memory) architectures. Having shown the feasibility of these approaches with regards to meeting medical image-to-mesh conversion requirements, the next step is to test both methods’ integration into a medical numerical simulation (while utilizing an error-based sizing function for PODM) to validate the end-user productivity of these approaches.
9 Acknowledgements
@@
This research was sponsored by the Richard T. Cheng Endowment at Old Dominion University, SURA grant No. C2024-FEMT-011-01, and the National Institute of General Medical Sciences of the National Institutes of Health under Award Number 1T32GM140911. This study was also partially supported by a pilot award from the Long Island Network for Clinical and Translational Science. The content is solely the authors’ responsibility and does not necessarily represent the official views of the National Institutes of Health. Additionally, this work was supported by the Anvil research computing cluster of Purdue University and the Wahab and Turing clusters at Old Dominion University. The authors would like to thank Dr. Fotis Drakopoulos and Dr. Christos Tsolakis for helping with the tools utilized in the adaptive pipeline of our approach. Additionally, we’d like to thank Dr. David Marcum for sharing the AFLR software utilized in this study, and Evangelia Frastali for helping to determine some optimal parameters for the pipeline methods tested in our evaluation. We’d also like to thank Spiros Tsalikis for discussions regarding the sizing function.